OpenStudy (anonymous):

In a three phase system, what is the instantaneos voltage of phase 2 when phase 1 is -85 volts and phase 3 is 116 volts?

OpenStudy (fifciol):

\[V(t)=V_0\cos(\omega t + \phi)\] phases are separated by 120 degrees ( 2pi/3) I can assume that theta is 0 at first phase so: \[-85=V_0\cos(\omega t )\] \[x=V_0\cos(\omega t + \frac{2 \pi}{3})\] \[116=V_0\cos(\omega t + \frac{4 \pi}{3})\] use also the equation: \[\cos(a+b)=cosa*cosb-sina*sinb\] and you should be able to find x :)

OpenStudy (fifciol):

or simply sum of instantaneous voltages of each phase must be zero: -85+x+116=0 x=-31